A 25 kg object moving with a velocity of 3.0 m/s to the right collides with a 15 kg object moving to the left at 6.0 m/s. Find the velocity of the 25 kg object after the collision, if the 15 kg object continues to move to the left but only at 0.30 m/s


1
Expert's answer
2021-11-09T16:35:59-0500

"m_1 = 25 \; kg \\

m_2 = 15 \; kg \\

u_1 = 3i \; m/s \\

u_2 = -6i \; m/s \\

v_2 = -0.3i \; m/s"

Momentum of the system before collision is

"p_i = m_1u_1 + m_2u_2 \\

= (25 \times 3)i + (-15 \times 6)i \\

= 75i -80i \\

= -5i \; kg \; m/s"

Momentum of the system after collision is

"p_f = m_1v_1 +m_2v_2 \\

= 25v_1 -(15 \times 0.3)i \\

= (25v -4.5i) \; kg \; m/s \\

p_f = p_i \\

25v -4.5i = -5i \\

25v = -0.5i \\

v = -0.02i"